# Changing the ranking in eigenvector centrality of a weighted graph by small perturbations

### v1

@@ Abstract.
In this article, we consider eigenvector centrality for the nodes of a graph and study the robustness (and stability) of this popular centrality measure. For a given weighted graph $\mathcal{G}$ (both directed and undirected), we consider the associated weighted adiacency matrix $A$, which by definition is a non-negative matrix. Eigenvector centrality consists of ranking the elements of the graph according to the corresponding entries of the Perron eigenvector of $A$, which is associated with the positive eigenvalue with largest modulus.
An indicator of the robustness of eigenvector centrality consists in looking for a nearby perturbed graph $\widetilde{\mathcal{G}}$, with the same structure as $\mathcal{G}$ (i.e., with the same vertices and edges), but with a weighted adiacency matrix $\widetilde{A}$ such that the highest $m$ entries ($m\geq 2$) of the Perron eigenvector of $\widetilde{A}$ coalesce, making the ranking at the highest level ambiguous. To compute a solution to this matrix nearness problem, a nested iterative algorithm is proposed that makes use of a constrained gradient system of matrix differential equations (possibly on a low-rank manifold) in the inner iteration and a one-dimensional optimization of the perturbation size in the outer iteration.
The proposed algorithm produces the optimal perturbation (i.e., the one with smallest Frobenius norm) of the graph, which causes the looked-for coalescence, which is a measure of the sensitivity of the graph. The methodology is formulated in terms of graphs but applies to any nonnegative matrix, with potential applications in fields like population models, consensus dynamics, economics, etc.
@@
Nicola Guglielmi acknowledges support from the Italian MUR within the PRIN 2022 Project “Advanced Numerical Methods for Time Dependent Parametric Partial Differential Equations with Applications”. He also thanks the support from the MIUR program for Departments of Excellence (project title of GSSI: Pattern analysis and engineering). Both authors acknowledge the MUR Pro3 joint project entitled “STANDS: Stability of Neural Dynamical Systems”. The authors are also affiliated to the INdAM-GNCS (Gruppo Nazionale di Calcolo Scientifico).
@@ 1. Introduction
A robustness indicator is useful in assessing the reliability of the measure. A very cheap and intuitively reasonable indicator of robustness consists of measuring the distance between the first and the $m$-th highest entries of the (suitably normalized) ranking eigenvector. However such a measure may be misleading, which means that even when this distance is small, this does not imply that the ranking is not robust, because - due to the structure of the graph - even a large modification of the weights of the graph might not be sufficient to annihilate the considered distance. For this reason we pose the problem in different terms and look for a closest graph $\widetilde{\mathcal{G}}$ to $\mathcal{G}$, which has in common with $\mathcal{G}$ the same sets of vertices and edges, but different weights, such that the largest entries of the Perron eigenvector coalesce. This problem translates into the search of a perturbed weighted adiacency matrix $\widetilde{A}$ to $A$ which preserves the sparsity pattern of $A$ and certain of its properties, and is characterized by colaescence of the first $m$-entries ($m\geq 2$) of the leading eigenvector.
@@
A main objective of this paper is to show how the robustness measure can be computed through the explicit computation of a nearby graph with ambiguous ranking. The proposed algorithm is an iterative algorithm, where in each step the leading eigenvector of the weighted adjacency matrix of a graph with perturbed weights is computed. For a large sparse graph (where the number of edges leaving any vertex is bounded independently of $n$), these computations can typically be done with a more favorable complexity in the number $n$ of vertices with respect to a general purpose eigensolver. A feature of this algorithm in common with recent algorithms for eigenvalue optimization (such as the ones described in [16, 18, 1]) is a two-level methodology for matrix nearness problems, where in an inner iteration a gradient flow drives perturbations of a fixed norm to the original matrix into a (local) minimum of a nonnegative functional that depends on eigenvalues and/or eigenvectors, while in an outer iteration the perturbation size is optimized such that the functional becomes zero.
The paper is organized as follows. In Section 2 we briefly mention recent related work by other authors working on similar centrality manipulation problems. In Section 3 we recall basic concepts of eigenvector centrality measures. In Subsection 3.1 we introduce the radius of robustness ($\varrho_{m}$) for eigenvector centrality as the distance to the closest graph with coalescent highest $m$ entries in the associated Perron eigenvector. In Subsection 3.2 we formulate a methodology to compute the $\varrho_{m}$ and the nearby ambiguous graph. This requires the solution of a structured matrix nearness problem. For this we propose a two-level iterative method, which is based on a gradient system of matrix differential equations in an inner iteration and a one-dimensional optimization of the perturbation size in an outer iteration, and we discuss algorithmic aspects. In Section 4 we provide the necessary theoretical background on first order perturbation theory of eigenvalues and eigenvectors. In Section 5 we describe the complete computational procedure to compute the radius of robustness. In Section 6 we present the results of numerical experiments where the robustness measure discussed here is computed for some examples of graphs. In Section 7 we briefly address the stability of the Perron eigenvector. Finally, in Section 8 we draw our conclusions.
@@ 3. Robustness and vulnerability of eigenvector centrality
be a weighted graph (either directed or undirected) with vertex set $\mathcal{V}=\{v_{1},\dots,v_{n}\}$, edge set $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$ and weighted adjacency matrix $A$. To each edge $(i,j)\in\mathcal{E}$ (from node $v_{j}$ to node $v_{i}$) is associated a positive weight $a_{ij}>0$; of course, for an undirected graph we have $a_{ji}=a_{ij}$ for all $i,j$. We set $a_{ij}=0$ for $(i,j)\notin\mathcal{E}$. Note that we exclude self-loops, hence $a_{ii}=0$ for all $i$. The weighted adjacency matrix of the graph is the non-negative matrix
@@
Eigenvector centrality (see, e.g., [32]) is based on the ranking of the entries of the suitably normalized Perron eigenvector $\mathrm{v}$.
@@
We show this exhibiting by the nearby graph $\widetilde{\mathcal{G}}$ (on the right picture) with weighted adiacency matrix $\widetilde{A}$ and corresponding Perron eigenvector $\widetilde{\mathrm{v}}$:
@@
Indeed, if a small perturbation is able to coalesce the highest $m$ entries of the eigenvector $\mathrm{v}$, it means that the ranking for the highest positions is not robust, especially in the presence of uncertainties, rounding errors, or any other form of noise in the weights; in general this is a symptom of the fact that the highest $m$ vertices are difficult to rank using eigenvector centrality and it may therefore be better A similar issue arises when a small perturbation makes the adjacency matrix $A$ reducible (i.e., the graph is no longer strongly connected), in which case the Perron eigenvector may no longer be defined, rendering eigenvector centrality no longer usable. Note that issue (II) above falls under this scenario.
In this paper we direct our attention mainly to issue (I), and we are interested in computing the smallest perturbation of $A$ such that the colaescence of the highest entries in the Perron vector occurs. To this end, in the next subsection we propose a sharp measure which is aimed at characterizing the reliability of eigenvector centrality. As a by-product, the knowledge of this perturbation allows the possibility of modifying the ranking by small changes to the weights and may be used for this purpose. In Section 7 we briefly discuss issue (II).
@@ 3.1. Robustness of the ranking and nearest ambiguous graph
$\varrho_{m}(\mathcal{G})=\>\inf_{\widetilde{\mathcal{G}}}{\rm dist}\left(\mathcal{G},\widetilde{\mathcal{G}}\right)=\inf_{\widetilde{A}}\|A-\widetilde{A}\|,$
@@
Note that we may choose to only perturb a subset of the edges, $\mathcal{E}^{\prime}\subset\mathcal{E}$, which would just change the constraint to $\widetilde{a}_{ij}\geq\delta$ for $(i,j)\in\mathcal{E}^{\prime}$. For example, we may choose to perturb only the edges which either are subject to uncertainties or to changes in time.
Indeed it is possible (although rare) that no graph $\widetilde{\mathcal{G}}$ exists which fulfils the constraints, in which case we say that the distance is $\infty$. Moreover, when there exist graphs fulfilling the constraints, we have that the infimum is indeed a minimum, by compactness.
@@ 3.2. Outline of the computational approach
the Perron eigenvector of $\widetilde{A}$ and define the mean of its highest $m$ entries as
@@
under the constraints of of preservation of the sparsity pattern of $A$, of unit Frobenius norm of $E$ and of non-negativity $A+\varepsilon E\geq 0$. This nonnegative functional measures the concentration of the highest $m$ entries of the Perron eigenvector $\mathrm{v}(A+\varepsilon E)$ with respect to their mean. Clearly the functional vanishes if and only if the $m$ highest entries of $\mathrm{v}(A+\varepsilon E)$ coalesce.
@@
  • Inner iteration: Given $\varepsilon>0$, we compute a real matrix $E=(e_{ij})\in\mathbb{R}^{n\times n}$ with the same sparsity pattern as $A$ (i.e., $e_{ij}=0$ if $a_{ij}=0$), with $E$ of unit Frobenius norm, with $(A+\varepsilon E)_{ij}\geq\delta$ (for $(i,j)\in\mathcal{E}$) such that the functional $F_{\varepsilon}(E)$ is minimized. The obtained minimizer is denoted by $E(\varepsilon)$.
@@ 4.1. First order perturbation theory for simple eigenvalues
The following standard perturbation result for eigenvalues is well-known; see e.g. [15, Theorem 1] and [27, Section II.1.1].
@@
Consider a continuously differentiable path of square real matrices $M(t)$ for $t$ in an open interval $I$. Let $\lambda(t)$, $t\in I$, be a continuous path of real simple eigenvalues of $M(t)$. Let $x(t)$ and $y(t)$ be left and right eigenvectors, respectively, of $M(t)$ to the eigenvalue $\lambda(t)$. Then, $x(t)^{\top}y(t)\neq 0$ for $t\in I$ and $\lambda$ is continuously differentiable on $I$ with the derivative (we denote ˙$\dot{\phantom{a}}=d/dt$)
@@ 4.2. First order perturbation theory for eigenvectors
For the derivative of eigenvectors we need the notion of group inverse (or reduced resolvent); see [31] as well as Kato ([27, Section I.5.3]).
@@
It is known from Meyer and Stewart ([31]) that if $M$ is normal, then its group inverse $M^{\#}$ is equal to the better known Moore–Penrose pseudoinverse $M^{\dagger}$. In general, the two pseudoinverses are not the same. They are, however, related by the following result, which is a special case of a more general result in [20, Appendix A] which is also verified directly.
@@
This result is used in the codes implemented by the authors for the computations involving the group inverse. Similar results are known in the Markov chain literature (see, e.g., [3]).
The group inverse appears in the following result on the derivative of eigenvectors, which is a variant of Theorem 2 of Meyer and Stewart ([31]).
@@ 5.1. Inner iteration
For a set of edges $\mathcal{E}$, we define $P_{\mathcal{E}}$ as the orthogonal projection from $\mathbb{R}^{n\times n}$ onto the vector subspace of all real matrices having the sparsity pattern determined by $\mathcal{E}$: for $A=(a_{ij})$,
$P_{\mathcal{E}}(A)\big{|}_{ij}:=\begin{cases}a_{ij}\,,&\text{if }(i,j)\in\mathcal{E}\,,\\ 0\,,&\text{otherwise.}\end{cases}$
For a fixed given weight matrix $A$ and for $\varepsilon>0$, we call a matrix $E=(e_{ij})\in\mathbb{R}^{n\times n}$ $\varepsilon$-feasible if
  • (ia)
    $E$ is of unit Frobenius norm.
  • (ib)
    $E=P_{\mathcal{E}}(E)$.
  • (ic)
    $E$ is symmetric (only for undirected graphs).
  • (ii)
    $(A+\varepsilon E)_{ij}\geq\delta$ for $(i,j)\in\mathcal{E}$.
@@
Similarly, if a few weights are related by a linear relationship, this can be easily included in the considered projection with respect to the Frobenius inner product.
5.2. Constrained gradient of the functional
Here we compute the gradient of the functional when restricted to matrices having the same sparsity pattern of $A$. The key idea is based on the fact that we are able to express the derivative of $F_{\varepsilon}(E)$ as $g(\dot{E})$ for a special function $g$. Using the simple fact that for vectors $b,c\in\mathbb{R}^{n}$ and a matrix $Z\in\mathbb{R}^{n\times n}$ one has

### v2

@@ Abstract.
In this article, we consider eigenvector centrality for the nodes of a graph and study the robustness (and stability) of this popular centrality measure. For a given weighted graph $\mathcal{G}$ (both directed and undirected), we consider the associated weighted adjacency matrix $A$, which by definition is a non-negative matrix. The eigenvector centralities of the nodes of $\mathcal{G}$ are the entries of the Perron eigenvector of $A$, which is the (positive) eigenvector associated with the eigenvalue with largest modulus. They provide a ranking of the nodes according to the corresponding centralities. An indicator of the robustness of eigenvector centrality consists in looking for a nearby perturbed graph $\widetilde{\mathcal{G}}$, with the same structure as $\mathcal{G}$ (i.e., with the same vertices and edges), but with a weighted adjacency matrix $\widetilde{A}$ such that the highest $m$ entries ($m\geq 2$) of the Perron eigenvector of $\widetilde{A}$ coalesce, making the ranking at the highest level ambiguous. To compute a solution to this matrix nearness problem, a nested iterative algorithm is proposed that makes use of a constrained gradient system of matrix differential equations in the inner iteration and a one-dimensional optimization of the perturbation size in the outer iteration.
The proposed algorithm produces the optimal perturbation (i.e., the one with smallest Frobenius norm) of the $A$ which causes the looked-for coalescence, which is a measure of the sensitivity of the graph. As a by-product of our methodology we derive an explicit formula for the gradient of a suitable functional corresponding to the minimization problem, leading to a characterization of the stationary points of the associated gradient system. This in turn gives us important properties of the solutions, revealing an underlying low rank structure. Our numerical experiments indicate that the proposed strategy outperforms more standard approaches based on algorithms for constrained optimization. The methodology is formulated in terms of graphs but applies to any nonnegative matrix, with potential applications in fields like population models, consensus dynamics, economics, etc.
@@
Nicola Guglielmi acknowledges support from the Italian MUR within the PRIN 2022 Project “Advanced Numerical Methods for Time Dependent Parametric Partial Differential Equations with Applications”. He also thanks the support from the MIUR program for Departments of Excellence (project title of GSSI: Pattern analysis and engineering).
Both authors acknowledge the MUR Pro3 joint project entitled “STANDS: Stability of Neural Dynamical Systems”. The authors are also affiliated to the INdAM-GNCS (Gruppo Nazionale di Calcolo Scientifico).
@@ 1. Introduction
A robustness indicator is useful in assessing the reliability of the measure. A very cheap and intuitively reasonable indicator of robustness consists of measuring the distance between the first and the $m$-th largest entries of the (suitably normalized) ranking eigenvector. However such a measure may be misleading, which means that even when this distance is small, this does not imply that the ranking is not robust, because - due to the structure of the graph - even a large modification of the weights of the graph might not be sufficient to annihilate the considered distance. For this reason we pose the problem in different terms and look for a closest graph $\widetilde{\mathcal{G}}$ to $\mathcal{G}$, which has in common with $\mathcal{G}$ the same sets of vertices and edges, but different weights, such that the largest entries of the Perron eigenvector coalesce. This problem translates into the search of a perturbed weighted adjacency matrix $\widetilde{A}$ to $A$ which preserves the sparsity pattern of $A$ and certain of its properties, and is characterized by coalescence of the first $m$ entries ($m\geq 2$) of the leading eigenvector.
@@
A main objective of this paper is to show how the robustness measure can be computed through the explicit computation of a nearby graph with ambiguous ranking. The proposed algorithm is an iterative algorithm, where in each step the leading eigenvector of the weighted adjacency matrix of a graph with perturbed weights is computed. For large sparse graphs with $n$ vertices and $O(n)$ edges, these computations can typically be done with a more favorable complexity in the number of vertices with respect to a general purpose eigensolver. A feature of this algorithm in common with recent algorithms for eigenvalue optimization (such as the ones described in [15, 16, 1]) is a two-level methodology for matrix nearness problems, where in an inner iteration a gradient flow drives perturbations of a fixed norm to the original matrix into a (local) minimum of a nonnegative functional that depends on eigenvalues and/or eigenvectors, while in an outer iteration the perturbation size is optimized such that the functional becomes zero.
The paper is organized as follows. In Section 2 we briefly mention recent related work by other authors working on similar centrality manipulation problems. In Section 3 we recall basic concepts of eigenvector centrality measures. In Subsection 3.1 we introduce the radius of robustness ($\varrho_{m}$) for eigenvector centrality as the distance to the closest graph with coalescent highest $m$ entries in the associated Perron eigenvector. In Subsection 3.3 we formulate a methodology to compute the $\varrho_{m}$ and the nearby ambiguous graph. This requires the solution of a structured matrix nearness problem. For this we propose a two-level iterative method, which is based on a gradient system of matrix differential equations in an inner iteration and a one-dimensional optimization of the perturbation size in an outer iteration, and we discuss algorithmic aspects. In Section 4 we provide the necessary theoretical background on first order perturbation theory of eigenvalues and eigenvectors. In Section 5 we describe the complete computational procedure to compute the radius of robustness. In Section 7 we present the results of numerical experiments where the robustness measure discussed here is computed for some examples of graphs. Finally, in Section 8 we draw our conclusions.
@@ 3. Robustness and vulnerability of eigenvector centrality
be a weighted graph (either directed or undirected) with vertex set $\mathcal{V}=\{v_{1},\dots,v_{n}\}$, edge set $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$ and weighted adjacency matrix $A$. To each edge $(i,j)\in\mathcal{E}$ (from node $v_{j}$ to node $v_{i}$: note the order) is associated a positive weight $a_{ij}>0$; of course, for an undirected graph we have $a_{ji}=a_{ij}$ for all $i,j$. We set $a_{ij}=0$ for $(i,j)\notin\mathcal{E}$. The weighted adjacency matrix of the graph is the non-negative matrix
@@
Eigenvector centrality (see, e.g., [28]) measures the importance of the nodes in $\mathcal{G}$ by the magnitude of the entries of the Perron eigenvector $\mathrm{v}$, allowing to rank the nodes. Note that the ranking is independent of the normalization used for $\mathrm{v}$.
@@
We show this exhibiting by the nearby graph $\widetilde{\mathcal{G}}$ (on the right picture) with weighted adjacency matrix $\widetilde{A}$ and corresponding Perron eigenvector $\widetilde{\mathrm{v}}$:
@@
Indeed, if a small perturbation is able to coalesce the highest $m$ entries of the eigenvector $\mathrm{v}$, it means that the ranking for the highest positions is not robust, especially in the presence of uncertainties, rounding errors, or any other form of noise in the weights; in general this is a symptom of the fact that the highest $m$ vertices are difficult to rank using eigenvector centrality and it may therefore be better to resort to different centrality measures. A similar issue arises when a small perturbation makes the adjacency matrix $A$ reducible (i.e., the graph is no longer strongly connected), in which case the Perron eigenvector may no longer be defined, rendering eigenvector centrality no longer usable. Note that issue (II) above falls under this scenario.
In this paper we direct our attention to issue (I), and we are interested in computing the smallest perturbation of $A$ such that the colaescence of the highest entries in the Perron vector occurs. To this end, in the next subsection we propose a sharp measure which is aimed at characterizing the reliability of eigenvector centrality. As a by-product, the knowledge of this perturbation allows the possibility of modifying the ranking by small changes to the weights and may be used for this purpose. Issue (II) is left as a topic for future work.
@@ 3.1. Robustness of the ranking and nearest ambiguous graph
$\varrho_{m}(\mathcal{G})=\>\inf_{\widetilde{\mathcal{G}}}{\rm dist}\left(\mathcal{G},\widetilde{\mathcal{G}}\right)\stackrel{{\scriptstyle\triangle}}{{=}}\inf_{\widetilde{A}}\|A-\widetilde{A}\|,$
@@
Note that we may choose to only perturb a subset of the edges, $\mathcal{E}^{\prime}\subset\mathcal{E}$, which would just change the constraint to $\widetilde{a}_{ij}\geq\delta$ for $(i,j)\in\mathcal{E}^{\prime}$. For example, we may choose to perturb only the edges which either are subject to uncertainties or to changes in time. Another possibility would be to perturb only the nodes adjacent to a set of control nodes as considered, e.g., in [29].
Indeed it is possible (although highly unlikely) that no graph $\widetilde{\mathcal{G}}$ exists which fulfills the constraints, in which case we say that the distance is $\infty$. Moreover, when there exist graphs fulfilling the constraints, we have that the infimum is indeed a minimum, by compactness.
3.2. The considered problem.
We are interested in the following problem of eigenvector-constrained minimization. We let
$\Sigma=\left\{B\in\mathbb{R}^{n\times n}\ \text{such that}\ b_{ij}\geq\delta\ \forall\ (i,j)\in\mathcal{E}\ \land\ b_{ij}=0\ \forall\ (i,j)\notin\mathcal{E}\right\}$
and (denoting with $\mathrm{u}$ the Perron eigenvector of $B\in\Sigma$)
(4) $\Omega=\left\{B\in\Sigma:\mathrm{u}_{i_{1}}=\cdots=\mathrm{u}_{i_{m}}\right\}$
where, for a given $m$, $i_{1},i_{2},\ldots,i_{m}$ are the indices of the $m$ largest entries of the Perron eigenvector $\mathrm{u}$ of $B$ (i.e., $B\mathrm{u}=\rho(B)\mathrm{u}$). The problem then reads
(5) $\min\|B-A\|_{F}\quad{\rm s.\,t.}\quad B\in\Omega\,.$
As we will see in the section on numerical experiments, the constraint defining the set $\Omega$ is challenging to deal with by means of classical optimization approaches.
In the next section we describe a reformulation of the problem which leads in natural way to a two-level approach, an inner eigenvector optimization problem and an outer root-finding one.
@@ 3.3. Outline of the computational approach
the (normalized) Perron eigenvector of $\widetilde{A}$ and define the mean of its largest $m$ entries as
@@
under the constraints of preservation of the sparsity pattern of $A$, of unit Frobenius norm of $E$ and of non-negativity $A+\varepsilon E\geq 0$. This nonnegative functional measures the dispersion of the largest $m$ entries of the Perron eigenvector $\mathrm{v}(A+\varepsilon E)$ with respect to their mean. Clearly the functional vanishes if and only if the $m$ highest entries of $\mathrm{v}(A+\varepsilon E)$ coalesce.
@@
  • Inner iteration: Given $\varepsilon>0$, we compute a smooth matrix flow $E(t)=(e_{ij}(t))\in\mathbb{R}^{n\times n}$, $t\geq 0$, with the same sparsity pattern as $A$ (i.e., $e_{ij}(t)=0$ if $a_{ij}=0$), with each $E(t)$ of unit Frobenius norm, with $(A+\varepsilon E(t))_{ij}\geq\delta$ (for $(i,j)\in\mathcal{E}$) and we minimize the functional $F_{\varepsilon}(E(t))$ over this matrix flow. The obtained minimizer is denoted by $E(\varepsilon)$.
@@
Two-level optimization
Due to the nested, or two-level form of problem (4)–(5), it is difficult to devise a single-level optimization approach to its solution.
The particular two-level methodology we propose has two main motivations. The first is that in this way we are able to approach the solution of the problem in a (norm) controlled way, which might be considered in analogy to a trust region method, allowing us – for example – to reach the solution by a sequence of matrices of smaller norm. The second is that in this way we obtain, as an important by-product, that the stationary points of the inner iteration have an explicit characterization in terms of the constrained gradient of the functional, therefore we are able to compute them explicitly by means of first-order eigenvector perturbation theory, as shown in the following Theorem 5.1. In particular, this reveals an underlying low-rank property that would otherwise remain hidden, and which can be exploited at the computational level.
For the solution of the inner iteration we propose a numerical integrator for the constrained gradient system associated to the functional. This is based on an explicit formula for the gradient, which we derive by using first-order eigenvector perturbation theory. Using alternative descent methods based on the exact knowledge of the gradient is also possible, and leads to similar results.
For the solution of the outer iteration (which aims to solve a root-finding problem) we consider a Newton-bisection method, which exploits an exact formula for the derivative of the function of which we aim to compute the smallest root.
@@ 4.1. First order perturbation theory for simple eigenvalues
The following standard perturbation result for eigenvalues is well-known; see e.g. [14, Theorem 1] and [23, Section II.1.1].
@@
Consider a continuously differentiable path of square real matrices $M(t)$ for $t$ in an open interval $I$. Let $\lambda(t)$, $t\in I$, be a continuous path of real simple eigenvalues of $M(t)$. Let $x(t)$ and $y(t)$ be left and right eigenvectors, respectively, of $M(t)$ corresponding to the eigenvalue $\lambda(t)$. Then, $x(t)^{\top}y(t)\neq 0$ for $t\in I$ and $\lambda$ is continuously differentiable on $I$ with the derivative (we denote ˙$\dot{\phantom{a}}=d/dt$)
@@ 4.2. First order perturbation theory for eigenvectors
For the derivative of eigenvectors we need the notion of group inverse (or reduced resolvent); see [27] as well as Kato ([23, Section I.5.3]).
@@
It is known from Meyer and Stewart ([27]) that if $M$ is normal, then its group inverse $M^{\#}$ is equal to the better known Moore–Penrose pseudoinverse $M^{\dagger}$. In general, the two pseudoinverses are not the same. They are, however, related by the following result, which is a special case of a more general result in [18, Appendix A] which is also verified directly.
@@
This result is used in the codes implemented by the authors for the computations involving the group inverse. Similar results are known in the Markov chain literature (see, e.g., [3]). The group inverse appears in the following result on the derivative of eigenvectors, which is a variant of Theorem 2 of Meyer and Stewart ([27]).
@@ 5.1. Inner iteration
For a set of edges $\mathcal{E}^{\prime}$, we define $P_{\mathcal{E}^{\prime}}$ as the orthogonal projection from $\mathbb{R}^{n\times n}$ onto the vector subspace of all real matrices having the sparsity pattern determined by $\mathcal{E}^{\prime}$: for $A=(a_{ij})$,
$P_{\mathcal{E}^{\prime}}(A)\big{|}_{ij}:=\begin{cases}a_{ij}\,,&\text{if }(i,j)\in\mathcal{E}^{\prime}\,,\\ 0\,,&\text{otherwise.}\end{cases}$
@@
Similarly, if a few weights are related by a linear relationship, this can be easily included in the considered projection with respect to the Frobenius inner product. It is sufficient to determine the projection onto the considered linear subspace and combine it with the one that takes into account the sparsity structure, which is also expressed as a linear subspace constraint. For details, see subsection 5.7.
From now on, unless otherwise specified, we shall consider $\mathcal{E}^{\prime}=\mathcal{E}$, that is, we perturb all edges of the graph.
Definition 5.1.
For a given weight matrix $A$ and $\varepsilon>0$, a matrix $E=(e_{ij})\in\mathbb{R}^{n\times n}$ is $\varepsilon$-feasible if
  • (a)
    $E$ is of unit Frobenius norm.
  • (b)
    $E=P_{\mathcal{E}}(E)$.
  • (c)
    $(A+\varepsilon E)_{ij}\geq\delta$ for $(i,j)\in\mathcal{E}$.
  • (d)
    $E$ is symmetric (only for undirected graphs).
5.2. Structured gradient of the functional (7)
Here we compute the gradient of the functional (7) when restricted to matrices having the same sparsity pattern of $A$. The key idea is based on the fact that we are able to express the derivative of $F_{\varepsilon}(E)$ as $g(\dot{E})$ for a special function $g$. Using the simple fact that for vectors $b,c\in\mathbb{R}^{n}$ and a matrix $Z\in\mathbb{R}^{n\times n}$ one has